in ap the 13th term is 22 and d is -1/2 then first term

Grade:10

1 Answers

Hari Charan reddy
3 years ago
In the given AP,
the first term, a=?
common difference, d= -1/2
and the 13th term = 22
we know that, the nth term of an AP = a + (n-1)d
so, 13th term = a+ 12d
then, a+12d = 22
         a+ 12(-1/2) = 22
         a – 6 = 22
therefore, a (first term) = 28.